# Causal Attention with Lookahead Keys

### v1

@@ 1 Introduction
Several recent works have proposed remedies to address these limitations on sentence embedding. BeLLM (Backward Dependency Enhanced LLM) [13] introduces backward dependencies which enables sentence embeddings to incorporate both past and future context, significantly improving performance on semantic textual similarity tasks and downstream applications. Echo Embeddings [21] duplicate the input and extracting representations from the second occurrence. The model effectively allows early tokens to attend to later ones and thus the model improves over classical LM embeddings without changing the architecture or requiring fine-tuning. Re-Reading (RE2) [26] enhances reasoning by prompting decoder-only LLMs to process the same question twice. The second pass provides tokens with access to global information from the first pass, capturing global dependencies and yielding stronger performance on arithmetic, commonsense, and symbolic reasoning tasks. Despite these advances, it remains unclear whether the above methods can benefit pretraining.
To address the shortcomings of standard causal attention in pretraining, we propose a novel attention mechanism, CAuSal aTtention with Lookahead kEys (CASTLE). In this approach, when generating the $(t+1)$-th token, we update keys of all preceding tokens so that keys of token $s$ ($1\leq s\leq t$) are able to additionally encode information from token $s+1$ to $t$. These keys are called lookahead keys. This design preserves the autoregressive structure while allowing the keys to evolve as the context unfolds. In Figure 1, we give an illustration of receptive fields of the keys in CASTLE. Although the mechanism appears recurrent, we establish a mathematical equivalence that avoids explicitly materializing the lookahead keys and enables efficient parallel training. We evaluate our approach on language modeling across multiple model scales. Experimental results show that CASTLE consistently outperforms standard causal attention in terms of validation perplexity and downstream task performance.
@@
The remainder of this paper is organized as follows. We discuss related work in Section 1.1. In Section 2.1, we elaborate on our motivations. In Section 2.2, we formally define CASTLE in its recurrent form. Section 2.3 proves an equivalent parallel formulation of CASTLE and develops efficient parallel training algorithms Section 2.4 introduces efficient inference algorithms together with the counterpart of the KV cache in CASTLE. Finally, Section 3 presents empirical results demonstrating the effectiveness of CASTLE across diverse tasks and model scales.
@@ 1.1 Related Work
Several studies have observed that the causal mask, by preventing tokens from accessing future information, can hinder a model’s ability to capture global context and thereby degrade its natural language understanding [13, 8, 21, 26, 30].
@@ 2.1 Motivations
$\begin{split}\text{CausalAttention}(\boldsymbol{\mathit{X}}^{L})=\text{row\_softmax}\Big{(}\frac{\boldsymbol{\mathit{Q}}\boldsymbol{\mathit{K}}^{\top}}{\sqrt{d}}+\boldsymbol{\mathit{M}}^{C}\Big{)}\boldsymbol{\mathit{V}}\in\mathbb{R}^{L\times d},\end{split}$ (1)
@@
To explain our motivations, we begin with the recurrent form of standard causal attention. Consider the case when generating the $(t+1)$-th token. Given contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$. Unless otherwise specified, all vectors in this paper are treated as row vectors rather than column vectors.
@@
$\begin{split}\text{causal-attention}(\boldsymbol{\mathit{X}}^{t})={\rm softmax}\left(\frac{\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{K}}_{t}}^{\top}}{\sqrt{d}}\right)\boldsymbol{\mathit{V}}_{t}=\frac{\sum_{s=1}^{t}\exp\big{(}\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big{)}\boldsymbol{\mathit{v}}_{s}}{\sum_{s=1}^{t}\exp\big{(}\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big{)}}\in\mathbb{R}^{1\times d}.\end{split}$
@@
This motivates us to propose a novel attention mechanism, causal attention with lookahead keys (CASTLE), i.e., when generating the $(t+1)$-th token, we first update keys $\boldsymbol{\mathit{k}}_{s}$ of preceding tokens $s$ with $s<t+1$ to additionally incorporate information from token $s+1$ to $t$. We refer to these as lookahead keys because their representations renew with the growing context. In this way, lookahead keys may lead to more accurate attention scores while preserving the autoregressive property of the model.
@@
We also remark that renewing the values $\boldsymbol{\mathit{v}}_{s}$ in a similar way to what we do for lookahead keys could potentially be beneficial. In this work, however, we focus on updating keys because doing so admits a mathematical equivalence (shown in Section 2.3) with $O(L^{2}d)$ computational complexity. Designing efficient parallel training algorithms that utilize lookahead values remains an interesting direction for future research.
@@ 2.2 Mathematical Definition in Recurrent Form
Throughout Section 2.2, we fix $t\in\left\{1,2,\cdots,L\right\}$ and consider the setting where $t$ tokens have already been generated and the model is about to generate the $(t+1)$-th token.
Denote the input contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$.
@@
For each preceding token $s$ ($1\leq s<t+1$), causal keys of token $s$ is a projection of $\boldsymbol{\mathit{x}}_{s}$, while lookaheads keys of token $s$ contain information from representations $\left\{\boldsymbol{\mathit{x}}_{s+1},\cdots,\boldsymbol{\mathit{x}}_{t}\right\}$. The receptive fields of causal keys and lookahead keys are illustrated in Figure 2.
@@
$\begin{split}\boldsymbol{\mathit{K}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{U},\boldsymbol{\mathit{V}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{U},\end{split}$
@@
$\begin{split}\boldsymbol{\mathit{K}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{C},\boldsymbol{\mathit{V}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{C}\end{split}$
@@
Causal Keys. The causal keys in CASTLE is defined similarly to the keys in standard causal attention. More specifically, causal keys are defined as
@@
The $s$-th row $\boldsymbol{\mathit{k}}_{s}$ of $\boldsymbol{\mathit{K}}^{C}_{t}$ satisfying $\boldsymbol{\mathit{k}}_{s}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{K}$ is the causal key of token $s$. Causal keys are static, i.e., the $s$-th rows of $\boldsymbol{\mathit{K}}^{C}_{t}$ and $\boldsymbol{\mathit{K}}^{C}_{t^{\prime}}$ are equal to each other whenever $t,t^{\prime}\geq s$.
Lookahead Keys. We utilize a structure similar to the attention mechanism to define lookahead keys.
@@
where $\boldsymbol{\mathit{M}}^{U}_{t}\in\mathbb{R}^{t\times t}$ is a mask matrix with $\boldsymbol{\mathit{M}}^{U}_{t}]_{ij}=0$ if $i<j$ and $\boldsymbol{\mathit{M}}^{U}_{t}]_{ij}=-\infty$ otherwise. The $s$-th row $\boldsymbol{\mathit{u}}^{t}_{s}$ of $\boldsymbol{\mathit{U}}^{t}$ is the lookahead key of token $s$. We remark that in $\boldsymbol{\mathit{u}}^{t}_{s}$, the superscript $t$ indicates that $\boldsymbol{\mathit{u}}^{t}_{s}$ is defined when $t$ tokens have already been generated and we are about to generate the $(t+1)$-th token, while the subscript $s$ indicates $\boldsymbol{\mathit{u}}^{t}_{s}$ is the $s$-th row of $\boldsymbol{\mathit{U}}^{t}$.
@@
An illustration for lookahead keys can be found in Figure 3. We have the following remarks regarding the definition of lookahead keys in (2).
@@
CASTLE in Recurrent Form. After defining causal keys and lookahead keys, we are ready to give the formula of CASTLE in recurrent form. To generate the $(t+1)$-th token, we utilize both the causal keys $\boldsymbol{\mathit{K}}^{C}_{t}\in\mathbb{R}^{t\times d}$ and the lookahead keys $\boldsymbol{\mathit{U}}^{t}$. More specifically, let the causal-key attention scores be
@@
where ${\rm SiLU}(\boldsymbol{\mathit{x}})=\boldsymbol{\mathit{x}}\odot{\rm sigmoid}(\boldsymbol{\mathit{x}})$.
Then, the output is calculated as
@@
We remark that $\rm SiLU$} is applied in (5) because our ablation study shows that $\boldsymbol{\mathit{p}}_{t}={\rm softmax}\left(\boldsymbol{\mathit{s}}^{C}_{t}-{\rm SiLU}\left(\boldsymbol{\mathit{s}}^{U}_{t}\right)\right)$ outperforms $\boldsymbol{\mathit{p}}_{t}={\rm softmax}\left(\boldsymbol{\mathit{s}}^{C}_{t}-\boldsymbol{\mathit{s}}^{U}_{t}\right)$ in downstream evaluation. We hypothesize that this improvement arises because many past tokens become “noise” as the context grows. The $\rm SiLU$} transformation effectively acts as a gate, regulating the degree to which each past token should be forgotten.
@@ 2.3 Efficient Parallel Training
In this section, we introduce our efficient parallel training algorithm. As discussed in Section 2.2, a straightforward materializing each $\boldsymbol{\mathit{U}}^{t}$ in parallel incurs at least $O(L^{3}d)$ computational costs. Such complexity makes training on large-scale datasets infeasible. To address this, we derive an equivalent parallel formulation of CASTLE that eliminates the need to materialize lookahead keys. This formulation then serves as the basis for designing an efficient algorithm.
@@
$\begin{split}\text{Attention}(\boldsymbol{\mathit{X}}^{L})=\begin{pmatrix}\text{attention}(\boldsymbol{\mathit{X}}^{1})\\ \text{attention}(\boldsymbol{\mathit{X}}^{2})\\ \vdots\\ \text{attention}(\boldsymbol{\mathit{X}}^{L})\end{pmatrix}\in\mathbb{R}^{L\times d_{\rm model}}.\end{split}$ (7)
The following theorem provides a parallel formulation of CASTLE that is equivalent to the recurrent form introduced in Section 2.2, but eliminates the need to explicitly materialize lookahead keys.
@@
Then, the outputs $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ as in (7) satisfies that
@@
Here, $\boldsymbol{\mathit{M}}^{C}$, $\widetilde{\boldsymbol{\mathit{M}}}^{C}$ are the causal masks which prevent tokens from attending to their future tokens, i.e., $\boldsymbol{\mathit{M}}^{C}_{ij}=0$ if $i\geq j$ and $\boldsymbol{\mathit{M}}^{C}_{ij}=$-$\infty$ otherwise; $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=1$ if $i\geq j$ and $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=0$ otherwise. $\boldsymbol{\mathit{M}}^{U}$ is the same as in (2), with $\boldsymbol{\mathit{M}}^{U}_{ij}=0$ if $i<j$ and $\boldsymbol{\mathit{M}}^{U}_{ij}=$-$\infty$ otherwise.
The proof of Theorem 1 is deferred to Appendix 5.
Efficient Parallel Training. Theorem 1 establishes the equivalence between the recurrent and parallel formulations of CASTLE. However, computing $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ directly from Theorem 1 still requires $\Omega(L^{3})$ complexity, since (8) involves matrix multiplications between $L\times L$ matrices.
@@
Given $\boldsymbol{\mathit{X}}^{L}$’s query, key and value matrices $\boldsymbol{\mathit{Q}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$, Algorithm 1 (forward pass) and Algorithm 3 (backward pass) enables efficient parallel training and can compute $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ and the gradients with computational complexity $O(L^{2}d)$ and space complexity $O(Ld)$.
The proof of Theorem 2 is given in Appendix 7.

### v2

@@ 1 Introduction
To address the shortcomings of standard causal attention in pretraining, we propose a novel attention mechanism, CAuSal aTtention with Lookahead kEys (CASTLE). In this approach, when generating the $(t+1)$-th token, we update keys of all preceding tokens so that keys of token $s$ ($1\leq s\leq t$) are able to additionally encode information from token $s+1$ to $t$. These keys are called lookahead keys. This design preserves the autoregressive structure while allowing the keys to evolve as the context unfolds. In Figure 1, we give an illustration of receptive fields of the keys in CASTLE. Although the mechanism appears recurrent, we establish a mathematical equivalence that avoids explicitly materializing the lookahead keys and enables efficient parallel training. We evaluate our approach on language modeling across multiple model scales. Experimental results show that CASTLE consistently outperforms standard causal attention in terms of validation perplexity and downstream task performance. We also introduce a variant, CASTLE-SWL, which applies sliding windows to lookahead keys. CASTLE-SWL has the same complexities with CASTLE in training and inference and preserves the performance gains of CASTLE while further improves efficiency in practice.
@@
The remainder of this paper is organized as follows. We discuss related work in Section 1.1. In Section 2.1, we elaborate on our motivations. In Section 2.2, we formally define CASTLE in its recurrent form. Section 2.3 proves an equivalent parallel formulation of CASTLE and develops efficient parallel training algorithms. Section 2.4 introduces efficient inference algorithms together with the counterpart of the KV cache in CASTLE. Finally, Section 3 presents empirical results demonstrating the effectiveness of CASTLE across diverse tasks and model scales.
@@ 1.1 Related Work
Several studies have observed that the causal mask, by preventing tokens from accessing future information, can hinder a model’s ability to capture global context and degrade its natural language understanding [15, 8, 23, 28, 33].
@@
PaTH Attention [30] is a novel positional encoding scheme which introduces data-dependent encodings based on cumulative products of Householder-like transformations. Each transformation is conditioned on the input, enabling the model to dynamically adapt positional information as the sequence progresses. PaTH is related to CASTLE in the sense that in inference formulation of PaTH, keys are also updated via a rank-1 update. However, both the update mechanism of keys and the parallel training formulation in PaTH differ substantially from those in CASTLE, and PaTH remains fundamentally a positional encoding, making it orthogonal to our approach.
Encoder-only Next Token Prediction (ENTP) [9] performs next-token prediction with encoder-only Transformers, where the keys are naturally re-computed at each position. It demonstrates stronger sample efficiency on small-scale language modeling and in-context learning benchmarks. However, the per-token compute in ENTP scales quadratically with sequence length and cubically over full sequences, which presents challenges for scaling.
@@ 2.1 Motivations
$\begin{split}\text{CausalAttention}(\boldsymbol{\mathit{X}}^{L})={\rm row\_softmax}\Big(\frac{\boldsymbol{\mathit{Q}}\boldsymbol{\mathit{K}}^{\top}}{\sqrt{d}}+\boldsymbol{\mathit{M}}^{C}\Big)\boldsymbol{\mathit{V}}\in\mathbb{R}^{L\times d},\end{split}$ (1)
@@
To explain our motivations, we begin with the recurrent form of standard causal attention. Consider the case when generating the $(t+1)$-th token. Given contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where the $s$-th row $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$. Unless otherwise specified, all vectors in this paper are treated as row vectors rather than column vectors.
@@
$\begin{split}\text{causal-attention}(\boldsymbol{\mathit{X}}^{t})={\rm softmax}\left(\frac{\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{K}}_{t}}^{\top}}{\sqrt{d}}\right)\boldsymbol{\mathit{V}}_{t}=\frac{\sum_{s=1}^{t}\exp\big(\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big)\boldsymbol{\mathit{v}}_{s}}{\sum_{s=1}^{t}\exp\big(\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big)}\in\mathbb{R}^{1\times d}.\end{split}$ (2)
@@
This motivates us to propose a novel attention mechanism, causal attention with lookahead keys (CASTLE), i.e., when generating the $(t+1)$-th token, we first update keys $\boldsymbol{\mathit{k}}_{s}$ of preceding tokens $s$ with $s<t+1$ to additionally incorporate information from token $s+1$ to $t$. We refer to these as lookahead keys because their representations renew with the growing context. In this way, lookahead keys may lead to more accurate attention scores while preserving the autoregressive property.
@@
We also remark that updating the values $\boldsymbol{\mathit{v}}_{s}$ similarly to lookahead keys could be beneficial while its efficient algorithm remains future study.
@@ 2.2 Mathematical Definition in Recurrent Form
Throughout Section 2.2, we fix $t\in\left\{1,2,\cdots,L\right\}$ and consider the setting where $t$ tokens have been generated and the model is generating the $(t+1)$-th token. Denote the input contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$.
@@
For each preceding token $s$ ($1\leq s<t+1$), causal keys of token $s$ is a projection of $\boldsymbol{\mathit{x}}_{s}$, while lookahead keys of token $s$ contain information from representations $\left\{\boldsymbol{\mathit{x}}_{s+1},\cdots,\boldsymbol{\mathit{x}}_{t}\right\}$. The receptive fields of causal keys and lookahead keys are illustrated in Figure 2.
@@
$\begin{split}\boldsymbol{\mathit{K}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{U}\in\mathbb{R}^{t\times d_{\rm model}},\ \boldsymbol{\mathit{V}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{U}\in\mathbb{R}^{t\times d_{\rm model}},\end{split}$
@@
$\begin{split}\boldsymbol{\mathit{K}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{C}\in\mathbb{R}^{t\times d_{\rm model}},\ \boldsymbol{\mathit{V}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{C}\in\mathbb{R}^{t\times d_{\rm model}}\end{split}$
@@
Causal Keys. The causal keys in CASTLE are defined similarly to the keys in standard causal attention. More specifically, causal keys are defined as
@@
The $s$-th row $\boldsymbol{\mathit{k}}_{s}$ of $\boldsymbol{\mathit{K}}^{C}_{t}$ satisfying $\boldsymbol{\mathit{k}}_{s}=\boldsymbol{\mathit{x}}_{s}\boldsymbol{\mathit{W}}^{C}_{K}$ is the causal key of token $s$. Causal keys are static, i.e., the $s$-th rows of $\boldsymbol{\mathit{K}}^{C}_{t}$ and $\boldsymbol{\mathit{K}}^{C}_{t^{\prime}}$ are equal to each other whenever $t,t^{\prime}\geq s$.
Lookahead Keys. We utilize a structure similar to the attention mechanism to define lookahead keys. An illustration for lookahead keys can be found in Figure 3.
@@
where $\boldsymbol{\mathit{M}}^{U}_{t}\in\mathbb{R}^{t\times t}$ is a mask matrix with
[…]
The $s$-th row $\boldsymbol{\mathit{u}}^{t}_{s}$ of $\boldsymbol{\mathit{U}}^{t}$ is the lookahead key of token $s$. We remark that in $\boldsymbol{\mathit{u}}^{t}_{s}$, the superscript $t$ indicates that $\boldsymbol{\mathit{u}}^{t}_{s}$ is defined when $t$ tokens have already been generated and we are about to generate the $(t+1)$-th token, while the subscript $s$ indicates $\boldsymbol{\mathit{u}}^{t}_{s}$ is the $s$-th row of $\boldsymbol{\mathit{U}}^{t}$.
@@
We have the following remarks regarding the definition of lookahead keys in (3).
@@
CASTLE in Recurrent Form. After defining causal keys and lookahead keys, we are ready to give the formula of CASTLE in recurrent form. To generate the $(t+1)$-th token, we utilize both the causal keys $\boldsymbol{\mathit{K}}^{C}_{t}\in\mathbb{R}^{t\times d}$ and the lookahead keys $\boldsymbol{\mathit{U}}^{t}$.
@@
More specifically, let the causal-key attention scores be
@@
where ${\rm SiLU}(\boldsymbol{\mathit{x}})=\boldsymbol{\mathit{x}}\odot{\rm sigmoid}(\boldsymbol{\mathit{x}})$. Then, the output is calculated as
@@
We remark that $\rm SiLU$} is applied in (7) because our ablation study indicates that it plays a crucial role in ensuring training stability. We hypothesize that this benefit arises because many past tokens become ‘noise’ as the context grows, and the SiLU transformation effectively acts as a gate, regulating the degree to which past tokens should be down-weighted.
@@
CASTLE-SWL in Recurrent Form. We propose a variant of CASTLE, termed CASTLE-SWL, where we apply Sliding Windows to Lookahead keys. More specifically, denote sliding window size by $W$. When generating the $(t+1)$-th token, lookahead keys of token $s$ ($s<t+1$) have access to information from $\left\{\boldsymbol{\mathit{x}}_{k}:s+1\leq k\leq\min\left\{s+W,t\right\}\right\}$. Formally, the definition of lookahead keys in CASTLE-SWL follows (3) with $\boldsymbol{\mathit{M}}^{U}_{t}$ defined as
[…]
An illustration for $\boldsymbol{\mathit{M}}^{U}_{t}$ applied to lookahead keys in CASTLE-SWL can be found in Figure 5.
Figure 5: An illustration of $\boldsymbol{\mathit{M}}^{U}_{t}$ in the lookahead keys of CASTLE-SWL, as defined in (9), when generating the $(t+1)$-th token. Each lookahead key of token $s$ ($s<t$) has access to representations $\left\{\boldsymbol{\mathit{x}}_{k}:s+1\leq k\leq\min\left\{t,,s+W\right\}\right\}$, i.e., up to $W$ subsequent tokens. Also, no information from tokens which are not yet generated is leaked and thereby preserving the autoregressive property. In this figure, we set $t=6$ and window size $W=3$. For example, lookahead keys of token 22 can “see” $\left\{\boldsymbol{\mathit{x}}_{3},\boldsymbol{\mathit{x}}_{4},\boldsymbol{\mathit{x}}_{5}\right\}$, while lookahead keys of token 44 can “see” $\left\{\boldsymbol{\mathit{x}}_{5},\boldsymbol{\mathit{x}}_{6}\right\}$. As analyzed in Section 2.3 and 2.4, CASTLE-SWL has the same complexities with CASTLE in both training and inference, however, it can further reduce FLOPs and improves efficiency in practice through the use of sliding windows.
Sliding window modifies only lookahead keys while causal keys remains unchanged. Apart from changing the definition of $\boldsymbol{\mathit{M}}^{U}_{t}$ from (4) to (9), all other components of CASTLE remain the same, yielding the definition of CASTLE-SWL.
The motivation for this design is that the semantic contribution of tokens may decay as the context unfolds. Allowing lookahead keys to aggregate information from distant tokens may introduce noise rather than useful signal.
In addition, although CASTLE-SWL has the same complexities with CASTLE in both training and inference as analyzed in Section 2.3 and 2.4, it can further reduce FLOPs and improves efficiency in practice through the use of sliding windows.
@@ 2.3 Efficient Parallel Training
In this section, we introduce our efficient parallel training algorithms. As discussed in Section 2.2, a straightforward materializing each $\boldsymbol{\mathit{U}}^{t}$ in parallel incurs at least $O(L^{3}d)$ computational costs. Such complexity makes training on large-scale datasets infeasible. To address this, we first derive an equivalent parallel formulation of CASTLE that avoids materializing lookahead keys, and then exploit the special structure of a matrix, which is expressible as a low-rank matrix multiplied by a mask, to reduce CASTLE’s training complexity to $O(L^{2}d)$. CASTLE-SWL naturally shares the same training complexity.
@@
$\begin{split}{\rm Attention}(\boldsymbol{\mathit{X}}^{L})=\begin{pmatrix}\text{attention}(\boldsymbol{\mathit{X}}^{1})\\ \text{attention}(\boldsymbol{\mathit{X}}^{2})\\ \vdots\\ \text{attention}(\boldsymbol{\mathit{X}}^{L})\end{pmatrix}\in\mathbb{R}^{L\times d_{\rm model}}.\end{split}$ (10)
The following theorem provides a unified parallel formulation for both CASTLE and CASTLE-SWL that is equivalent to their recurrent forms introduced in Section 2.2. This formulation then serves as the basis for designing an efficient algorithm. Its proof is in Appendix 6.
@@
Then, the outputs ${\rm Attention}(\boldsymbol{\mathit{X}}^{L})$ as in (10) satisfies that
@@
Here, $\boldsymbol{\mathit{M}}^{C}$, $\widetilde{\boldsymbol{\mathit{M}}}^{C}$ are the causal masks which prevent tokens from attending to their future tokens, i.e., $\boldsymbol{\mathit{M}}^{C}_{ij}=0$ if $i\geq j$ and $\boldsymbol{\mathit{M}}^{C}_{ij}=$-$\infty$ otherwise; $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=1$ if $i\geq j$ and $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=0$ otherwise. For CASTLE, $\boldsymbol{\mathit{M}}^{U}=\boldsymbol{\mathit{M}}^{U}_{L}\in\mathbb{R}^{L\times L}$ with $\boldsymbol{\mathit{M}}^{U}_{L}$ defined in (4), while for CASTLE-SWL, $\boldsymbol{\mathit{M}}^{U}=\boldsymbol{\mathit{M}}^{U}_{L}\in\mathbb{R}^{L\times L}$ with $\boldsymbol{\mathit{M}}^{U}_{L}$ defined in (9).
Efficient Parallel Training. Theorem 1 establishes the equivalence between the recurrent and parallel formulations for CASTLE and CASTLE-SWL. However, computing ${\rm Attention}(\boldsymbol{\mathit{X}}^{L})$ directly from Theorem 1 still requires $\Omega(L^{3})$ complexity, since (11) involves matrix multiplications between $L$-by-$L$ matrices.
@@
Given $\boldsymbol{\mathit{X}}^{L}$’s query, key and value matrices $\boldsymbol{\mathit{Q}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$, for both CASTLE and CASTLE-SWL, Algorithm 1 (forward pass) and Algorithm 3 (backward pass) enable efficient parallel training and can compute ${\rm Attention}(\boldsymbol{\mathit{X}}^{L})$ and the gradients with computational complexity $O(L^{2}d)$ and space complexity $O(Ld)$.